OpenStudy (anonymous):

help ? before my time expires Solve the equation for y and then find the value of y for the given x-value. 5x + y = 10

OpenStudy (anonymous):

@Phebe

OpenStudy (aum):

5x + y = 10 To solve for y, isolate y. subtract 5x from both sides: y = 10 - 5x "find the value of y for the given x-value." What is the x-value? You have not provided the x-value.

OpenStudy (phebe):

x = 2-y/5

OpenStudy (phebe):

@aum can u keep explaing it

OpenStudy (aum):

@Phebe Yes, I can explain it. The problem asks to solve for y and not solve for x as done in your first reply.

OpenStudy (phebe):

ohhhhhhhhhhhh thats sooo trueeee

OpenStudy (phebe):

sorry

OpenStudy (aum):

no big deal. @cathrineT Part of the problem is missing. You have not provided the x-value.

OpenStudy (aum):

The problem says: " find the value of y for the given x-value." We need the x-value.

OpenStudy (anonymous):

I didnt mean to give you the x-value they just want the equation for y I already have x

OpenStudy (aum):

y = 10 - 5x

OpenStudy (anonymous):

I have to solve it ?

OpenStudy (anonymous):

@aum